OpenStudy (adrianna.gongora):

Solve x2 + 15x = -56. Options x = 8 and x = 7 x = -8 and x = -7 x = -8 and x = 7 x = 8 and x = -7

Directrix (directrix):

x2 + 15x = -56 x^2 + 15x + 56 = 0 Factor. @Adrianna.Gongora

OpenStudy (mayamartinez):

im going to do this

OpenStudy (mayamartinez):

it looks fun

OpenStudy (adrianna.gongora):

It's either A or B ;-; I know that but I don't know which one..

OpenStudy (adrianna.gongora):

@Directrix

Directrix (directrix):

What are two numbers that multiply to 56 AND also add to 15? @Adrianna.Gongora

OpenStudy (adrianna.gongora):

7 x 8 = 56 7 + (8) = 15

OpenStudy (adrianna.gongora):

It's A I think

Directrix (directrix):

x^2 + 15x + 56 = 0 (x + 8) * (x + 7) = 0 x + 8 = 0 or x + 7 = 0 x = ? or x = ?

OpenStudy (adrianna.gongora):

(X + 7 ) (X + 8) soo X=7 and X=8

Directrix (directrix):

No.

OpenStudy (adrianna.gongora):

ughh ;-;

Directrix (directrix):

x + 8 = 0 or x + 7 = 0 x + 8 = 0 or x + 7 = 0 x = 0 - 8 x = 0 - 7 x = -8 x = -7

OpenStudy (adrianna.gongora):

;-; ooh my lord so it's b?

Directrix (directrix):

x = -8 and x = -7 Correct.

OpenStudy (adrianna.gongora):

That's what I was thinking but then you put (x + 8) * (x + 7) = 0 and I didn't see negatives so I assumes positive ._. I get messed up with negatives and positives sometimes ;-; Thanks though

Directrix (directrix):

I should have posted that I was applying the Zero Product Property. See attachment and let me know if that clears the confusion.

OpenStudy (adrianna.gongora):

That actually makes a lot of sense :) Thank you !

Directrix (directrix):

You are welcome.